OpenStudy (anonymous):

Given the function g(x) = 8x − 2, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values. The value of g(−2) is larger than the value of g(4). The value of g(−2) is the same as the value of g(4). The value of g(−2) is smaller than the value of g(4). The values of g(−2) and g(4) cannot be compared.

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

Plug in -2 and 4 where x is in g(x)

OpenStudy (anonymous):

g(x) = 8x - 2 x = -2 Therefore g(-2) = 8(-2) - 2 g(-2) = -16 - 2 = -18 Do the same for 4

OpenStudy (anonymous):

8(4)-2

OpenStudy (anonymous):

Okay, then what?

OpenStudy (anonymous):

g(4) = 8(4) - 2 g(4) = 32 - 2

OpenStudy (anonymous):

30

hartnn (hartnn):

since g(x) has positive slope, g(4) will be larger than g(-2) or u can solve and see it like that^^

OpenStudy (anonymous):

right, so compare and contrasting, is 30 greater, lesser or equal to -18?

OpenStudy (anonymous):

Arielllllllllll just help u, i think THEY ARE RIGHT

OpenStudy (anonymous):

umm' noo

hartnn (hartnn):

why no? as we can see f(4)=30 is > f(-2)=-18 so option C is correct choice.

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

and 1 more

OpenStudy (anonymous):

If A = {black, pink} and B = {red, orange, yellow, green, blue, indigo, violet}, what is A B? {black, pink} {red, orange, yellow, green, blue, indigo, violet, black, pink} {all colors except black and pink} { }

hartnn (hartnn):

what is in between A and B ??

OpenStudy (anonymous):

umm cofuse

hartnn (hartnn):

i mean in the question what is A B? there must be a sign between them

hartnn (hartnn):

\[\cup ??\]

OpenStudy (anonymous):

union

hartnn (hartnn):

then AUB contains 9 elements, which option has 9 elements??

OpenStudy (anonymous):

second

hartnn (hartnn):

thats it, 2nd option it is